My son (14) says there is no solution to his homework question, is he correct

(w+x)(y+z)
w*y=4
x*z= -15
(w*z)+(x*Y)=1

Answer by stanbon(75887)

My son (14) says there is no solution to his homework question, is he correct
(w+x)(y+z)
w*y=4
x*z= -15
(w*z)+(x*y)=1
----------------
====================
(w+x)(y+z) = wy+wz+xy+xz
---
Substitute to get:
4 + 1 -15 = -10
=========================
Cheers,
Stan H.